OpenStudy (anonymous):

Evaluate: 4 (5 + 2 x 7) -3^3

OpenStudy (mathstudent55):

First work inside the parentheses. Inside the parentheses there is a multiplication and an addition. In the order of operations, the multiplication comes first. Do the multiplication inside the parentheses and copy down everything else exactly as it is. What do you get?

OpenStudy (anonymous):

Thank you! It's not that I didn't know how to do the problem, my issue was remembering whether or not to do the multiplication inside the parentheses first. Thank you, again!

OpenStudy (mathstudent55):

What did you get as a final answer?

OpenStudy (anonymous):

I got 67.

OpenStudy (mathstudent55):

That's not what I got. Let's do it together one step at a time. It's important to do the operations in the correct order, and it is also important to copy down everything else at every step so you don't forget any number or operation.

OpenStudy (mathstudent55):

4 (5 + 2 x 7) -3^3 = 4(5 + 14) - 3^3

OpenStudy (mathstudent55):

That was the multiplication inside the parentheses. Now we do the addition inside the parentheses.

OpenStudy (anonymous):

That would be 19. I did get that far in the equation.

OpenStudy (mathstudent55):

4 (5 + 2 x 7) -3^3 = 4(5 + 14) - 3^3 = 4(19) - 3^3 Now we do the exponent. Notice the exponent is 3, not 2. 3^3 = 27. 3^2 = 9. We need 3^3 = 27. = 4(19) - 27 Now we do the multiplication: = 76 - 27 Now we do the subtraction = 49

OpenStudy (mathstudent55):

I think you may have just done 3^2 = 9 instead of 3^3 = 27.

OpenStudy (anonymous):

3 x 3 is nine.

OpenStudy (anonymous):

Oh! Never mind! Whoops!

OpenStudy (anonymous):

You're right. I'm wrong.

OpenStudy (mathstudent55):

Wait. Be careful with exponents.

OpenStudy (anonymous):

Thank you so much! I do have it now. I'm going over review and it's harder this year than it was last.

OpenStudy (mathstudent55):

Multiplication is 3 x 3 = 9. If multiplication already exists, and 3 x 3 = 9, why would mathematicians invent exponents for them to mean the same? That means 3^3 does not mean 3 * 3. Exponents exist to show a series of multiplications in a very short way of writing it. It's true that 3^2 and 3*3 take as many characters to wrIte, but now look at 3^7. That is a very short way of writing 3 * 3 * 3 * 3 * 3 * 3 * 3, which, BTW, equals 2187.

OpenStudy (mathstudent55):

Ok, you're welcome. Keep studying and asking questions. You'll do well.

OpenStudy (anonymous):

Would you mind helping me with another? Gosh, I honestly hate Algebra. It's actually simple, but messes me up.

OpenStudy (mathstudent55):

No problem. Pls start a new post. I'll help you there.

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

wlcm